OpenStudy (anonymous):

Given a triangle with b = 4, c = 10, and A = 84º, what is the length of a? Round to the nearest tenth. Find the area of the triangle with a = 11.8, b = 12.6, c = 14.8. Round to the nearest tenth.

OpenStudy (anonymous):

(1) This is based on Law of Cosines: \[c^2=a^2+b^2-2ab \cos(C)\] In this case, we want a, so let's setup the law of cosines to solve for a: \[a^2=b^2+c^2-2bc \cos(A)\] Substituting in values we have: \[a=\sqrt{(4)^2+(10)^2-2(4)(10)\cos(84)}\]\[a=\sqrt{116-80\cos(84)}\] Therefore: \[a=10.4\] (2) This is based on Heron's Formula: \[Area=\sqrt{p(p-a)(p-b)(p-c)}\]Where: \[p=\frac{ a+b+c }{ 2 }\] Rewriting Heron's Formula gives us: \[Area=\sqrt{(\frac{ a+b+c }{ 2 })(\frac{ a+b+c }{ 2 }-a)(\frac{ a+b+c }{ 2 }-b)(\frac{ a+b+c }{ 2 }-c)}\] Solving a+b+c / 2 we get: \[p=\frac{ a+b+c }{ 2 }=\frac{ 11.8+12.6+14.8 }{ 2 }=19.6\] Substituting in values we have: \[Area= \sqrt{(19.6)(19.6-11.8)(19.6-12.6)(19.6-14.8)}\] Therefore: \[Area = 71.7\]